# Minibatch Optimal Transport and Perplexity Bound Estimation in Discrete Flow Matching

### v3

@@ Abstract
Discrete flow matching, a recent framework for modeling categorical data, has shown competitive performance with autoregressive models. However, unlike continuous flow matching, the rectification strategy cannot be applied due to the stochasticity of discrete paths, necessitating alternative methods to minimize state transitions. We propose a dynamic-optimal-transport-like minimization objective and derive its Kantorovich formulation for discrete flows with convex interpolants, where transport cost depends solely on inter-state similarity and can be optimized via minibatch strategies. In the case of bag-of-words (BoW) sourced flows, we show that such methods can reduce the number of transitions up to 8 times (1024 to 128) to reach the same generative perplexity without compromising diversity. Additionally, path nondeterminism in discrete flows precludes an instantaneous change-of-variables analogue, preventing precise probability estimation available to continuous flows. We therefore propose two upper bounds on perplexity, enabling principled training, evaluation and model comparison. Finally, we introduce Multimask Flows which outperform masked flows in generative perplexity, particularly when utilizing minibatch Optimal Transport, without sacrificing diversity.
@@ 1 Introduction
Modeling data distributions is central to machine learning. For continuous data, diffusion and flow models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020b; Lipman et al., 2023) have shown impressive results in generation and density estimation (Song et al., 2021a; c; Chen et al., 2018). Rectified flows particularly excel by enabling high-quality generation with few integration steps. However, these continuous models lag behind autoregressive models on categorical data (Chen et al., 2023; Gulrajani & Hashimoto, 2024; Li et al., 2022; Dieleman et al., 2022; Strudel et al., 2022).
@@
To address this, recent work has developed discrete diffusion (Austin et al., 2021; Campbell et al., 2022; Meng et al., 2022; Lou et al., 2024; Sahoo et al., 2024; Ou et al., 2024; Shi et al., 2024) and discrete flow models (Campbell et al., 2024; Gat et al., 2024), better suited for categorical data. These models can accelerate generation and, unlike autoregressive models, naturally enable infilling. We focus on discrete flow matching (DFM), which expands the design space beyond discrete diffusion by allowing arbitrary couplings and inner dynamics. While discrete and continuous flows share similarities that facilitate adapting continuous flow matching results, fundamental differences remain. These arise primarily from the nonexistence of a DFM formulation with deterministic sample paths.
@@
A major implication of non-deterministic sample paths is that we cannot use the rectification strategy from Liu et al. (2022). Since paths in discrete flows are sequences of states, we explore minimizing the number of jumps between states, which can be interpreted as the discrete analogue of path length minimization. Using similarity measures between states, we minimize jumps weighted by dissimilarity, yielding a weighted path-length-oriented dynamic formulation of optimal transport (OT) for discrete flow matching. We derive its Kantorovich formulation, where the cost function depends only on the similarity measure and can be optimized using minibatch strategies (Tong et al., 2024; Fatras et al., 2021). This gives a categorical Benamou-Brenier-type theorem when conditional flows are convex interpolants, i.e., the categorical equivalent of shortest-path continuous flows. When the similarity measure in the dynamic formulation is the discrete metric, the cost function in the Kantorovich formulation becomes the Hamming distance. When the similarity measure is the $L_{2}$ norm, the cost function is also $L_{2}$, mirroring the continuous case.
@@
An additional implication of stochastic crossing paths in DFM is that we cannot use an equivalent of the instantaneous change of variable formula (Chen et al., 2018) for probability estimation. Thus, other approaches are needed for estimating the perplexity. Inspired by bounds in Lou et al. (2024); Haxholli et al. (2025), we derive two upper bounds on perplexity for discrete flow matching. These bounds enable theoretically grounded training, model evaluation and comparison with other methods.
@@
Experiments show that minibatch-OT significantly reduces jumps in small-scale experiments and, in realistic settings (GPT2-sized model on OWT), reduces inference steps up to 8-fold (from 1024 to 128) to achieve the same generative perplexity. We also introduce multimask flow (DFM-MM), which outperforms masked DFM in terms of generative perplexity without sacrificing diversity, in particular when combined with minibatch-OT. Finally, we demonstrate that our derived bounds enable comparisons with autoregressive and discrete diffusion models.
@@
  • We show minibatch OT reduces inference steps up to 8-fold (1024 to 128) while maintaining generative perplexity on GPT2-scale models. Finally, we introduce multimask flow (DFM-MM), which surpasses masked DFM models in generative perplexity without compromising diversity, with further gains achieved when applying OT.
@@ 2 Preliminaries and Notation
A summary of Discrete Flow Matching is provided below. While the following preliminary is self-contained, we also provide an introduction to the discrete diffusion framework in Appendix D.
@@ 2.1.2 Conditional Probability Flows
$\text{where }p_{t}^{i}(x^{i}|x_{0},x_{1})=(1-k_{t})\delta_{x_{0}^{i}}(x^{i})+k_{t}\delta_{x_{1}^{i}}(x^{i}),\text{with }k_{0}=0,k_{1}=1\text{ and increasing }k_{t}.$ (4)
It is clear that this definition of $p_{t}(x|x_{0},x_{1})$ satisfies the conditions $p_{0}(x|x_{0},x_{1})=\delta_{x_{0}}(x)$ and $p_{1}(x|x_{0},x_{1})=\delta_{x_{1}}(x)$. In addition, Gat et al. (2024) show that component $i$ of the conditional probability velocity $u_{t}(x,z|x_{0},x_{1})$ corresponding to the flow defined in Equations (3) and (4) is
@@
Furthermore, they show that the probability velocity corresponding to the unconditional flow $p_{t}(z)$ can be written as
$u_{t}^{i}(x^{i},z)=\sum_{x_{0},x_{1}\in\mathcal{D}}u_{t}^{i}(x^{i},z|x_{0},x_{1})p(x_{0},x_{1}|z)dx_{0}dx_{1},$ (6)
which in the case of Equations (4) and (5) implies, $u_{t}^{i}(x^{i},z)=\frac{\dot{k}_{t}}{1-k_{t}}\left[p_{1|t}^{i}(x^{i}|z)-\delta_{z}(x^{i})\right].$ One then approximates $u_{t}^{i}(x^{i},z)$ by simply modeling $p^{i}_{1|t}(x^{i}|z)$ with a neural network $p^{i}_{1|t}(x^{i}|z;\theta)$ using the cross entropy loss $L$,
$\mathbb{E}_{t\sim U(0,1)}\mathbb{E}_{x_{0},x_{1}\sim\pi(x_{0},x_{1})}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0},x_{1})}\sum_{i=1}^{L}\log p^{i}_{1|t}(x_{1}^{i}|x_{t};\theta).$ (7)
It should be mentioned that in Gat et al. (2024), the definition of $p_{t}^{i}(x^{i}|x_{0},x_{1})$ is given in a more general form, but here we focus on this specific case for the sake of simplicity and since this formulation corresponds to shortest path conditional flows in the continuous framework, that is $X_{t}=(1-t)X_{0}+tX_{1}$.
@@ 3 Transition Reduction Objectives in Discrete Flow Matching
We first introduce our dynamic optimal transport formulation for discrete flows with per-position convex interpolants and derive its equivalent static (Kantorovich) formulation. We then explain the novelty of our results in this particular flow setting. Finally, since masked flows (Gat et al., 2024) achieve the best practical performance but cannot leverage OT, we introduce multimasked flows. The latter maintain the performance and time-independent predictive probabilities of masked flows while enabling the application of OT approaches.
@@ 3.1 Dynamic OT Formulation of Discrete Flows with convex interpolants and its Static Equivalent
with endpoints fixed at the source and target distributions. This objective is closely related to minimizing expected path length, but the squared speed penalizes large velocities more strongly. Moreover, by the Benamou–Brenier theorem (Benamou & Brenier, 2000; Tong et al., 2024), the infimum of Equation (8) equals the infimum of the Kantorovich transport with quadratic cost,
$\int c(x_{0},x_{1})\pi(x_{0},x_{1})dx_{0}dx_{1},\text{ where }c(x_{0},x_{1})=\|x_{0}-x_{1}\|^{2},$ (9)
@@
$\int_{0}^{1}\sum_{x_{t}}\frac{1}{2}p(x_{t})\left[\sum_{i=1}^{L}\left(\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})-u_{t}^{i}(x_{t}^{i},x_{t})\right)\right]dt=\int_{0}^{1}\sum_{x_{t}}p(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})dt,$ (10)
@@
For these reasons, we define the categorical dynamic objective more generally as follows:
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\left[\sum_{i=1}^{L}\sum_{x^{i}}u_{t}^{i}(x^{i},x_{t})s(x^{i},x_{t}^{i})\right]dt=\int_{0}^{1}\sum_{x_{t}}p(x_{t})\left[\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})s(x^{i},x_{t}^{i})\right]dt,$ (11)
@@
Let $\pi(x_{0},x_{1})$ be the joint distribution of $x_{0}$ and $x_{1}$, and let $p_{t}$ be a flow defined as in Equations (2, 3, 4) that transforms $p=\int\pi(x_{0},x_{1})dx_{1}$ into $q=\int\pi(x_{0},x_{1})dx_{0}$. In this setting, the dynamic formulation given in Equation (11) equals the Kantorovich formulation:
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})s(x^{i},x_{t}^{i})dt=\sum_{x_{0},x_{1}}c(x_{0},x_{1})\pi(x_{0},x_{1}),$ (12)
@@
Theorem 3.1 focuses on a particular flow setting which covers all of the usual flows encountered in the literature: per-position discrete flows on sequence space, specifically with convex interpolants. This structured setting enables results for a broader class of cost functions than the classical Benamou–Brenier theorem. Within this framework, we introduce a general similarity function $s(\cdot,\cdot)$ and define our dynamic objective using path length rather than kinetic energy; that is, discrete jump rates and weighted jump counts instead of quadratic velocity fields. This naturally accommodates more general discrepancy measures beyond $\|\cdot\|^{2}$. This restriction proves advantageous, as our Benamou–Brenier-type theorem holds for arbitrary similarity functions $s$, not just squared Euclidean distances. Further, since the flow dynamics are fully determined by the coupling $\pi$, we optimize only over couplings, not entire flows. For each $\pi$, the induced convex-interpolant flow yields a dynamic loss that exactly equals the static Kantorovich loss for the same $\pi$. Taking the infimum over all couplings is therefore purely formal, since equality holds pointwise before optimization.
@@ 3.2 Multi-mask Flows
Standard masked DFM (Gat et al., 2024) uses a Dirac distribution at the fully masked sequence as its source, admitting only the trivial coupling. To enable meaningful couplings, we introduce multimask flow (DFM-MMLM) with $V_{s}=50,257$ special mask tokens $m_{1},m_{2},...,m_{V_{s}}$}, all distinct from data tokens. Source sequences are uniformly sampled combinations of these masks. Our total vocabulary thus contains $V=V_{s}+V_{d}$ tokens: $V_{d}=50,257$ data tokens $x_{1},...,x_{V_{d}}$} from our tokenizer and $V_{s}=50,257$ mask tokens. At $t=0$, only mask tokens appear while data tokens are assigned zero probability. This design provides two advantages: denoising probabilities remain time-independent as in masked flow (see Appendix A.6), and mask embeddings are fully unrestricted since they are untied from data-token embeddings. Effectively, we create a "fictitious grid" where each $L$-length sequence carries mass $\frac{1}{50257^{L}}$. The flow transports this mass to the data grid, enabling minibatch OT.
@@ 4 Upper Bounds on the Perplexity in Discrete Flow Matching
Perplexity is a key metric for language models, making it essential to calculate or bound it in the DFM framework. While Appendix A.5 provides a precise but computationally intractable formula, the next two subsections present practical bounds. These bounds serve as both principled training objectives and effective evaluation metrics, offering intrinsic and objective assessment.

### v4

@@ Abstract
Discrete flow matching, a recent framework for modeling categorical data, has shown competitive performance with autoregressive models. However, unlike continuous flow matching, the rectification strategy cannot be applied due to the stochasticity of discrete paths, necessitating alternative methods to minimize state transitions. We propose a dynamic-optimal-transport-like minimization objective and derive its Kantorovich formulation for discrete flows with convex interpolants, where transport cost depends solely on inter-state similarity and can be optimized via minibatch strategies. We show that such methods can reduce the number of transitions up to 32 times (1024 to 32) to reach the same generative perplexity without compromising diversity. Additionally, path nondeterminism in discrete flows precludes an instantaneous change-of-variables analogue, preventing precise probability estimation available to continuous flows. We therefore propose two upper bounds on perplexity, enabling principled training, evaluation and model comparison. Finally, we introduce Multimask Flows which outperform masked flows in generative perplexity without compromising diversity, particularly when utilizing minibatch Optimal Transport.
Machine Learning, ICML
@@ 1 Introduction
Modeling data distributions is central to machine learning. For continuous data, diffusion and flow models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020b; Lipman et al., 2023) have shown impressive results in generation and density estimation (Song et al., 2021a, c; Chen et al., 2018). Rectified flows particularly excel by enabling high-quality generation with few integration steps.
However, these continuous models lag behind autoregressive models on categorical data (Chen et al., 2023; Gulrajani and Hashimoto, 2024; Li et al., 2022; Dieleman et al., 2022; Strudel et al., 2022).
@@
To address this, recent work has developed discrete diffusion (Austin et al., 2021; Campbell et al., 2022; Meng et al., 2022; Lou et al., 2024; Sahoo et al., 2024; Ou et al., 2024; Shi et al., 2024) and discrete flow models (Campbell et al., 2024; Gat et al., 2024), better suited for categorical data. Such models can accelerate generation and, unlike autoregressive models, naturally enable infilling. We focus on discrete flow matching (DFM), which expands the design space beyond discrete diffusion by allowing arbitrary couplings and inner dynamics. While discrete and continuous flows share similarities that facilitate adapting continuous flow matching results, fundamental differences remain. These arise primarily from the nonexistence of a DFM formulation with deterministic sample paths.
@@
A major implication of non-deterministic sample paths is that we cannot use the rectification strategy from Liu et al. (2022). Since paths in discrete flows are sequences of states, we explore minimizing the number of jumps between states, which can be interpreted as the discrete analogue of path length minimization. Using similarity measures between states, we minimize jumps weighted by dissimilarity, yielding a weighted path-length-oriented dynamic formulation of optimal transport (OT) for discrete flow matching. We derive its Kantorovich formulation, where the cost function depends only on the similarity measure and can be optimized using minibatch strategies (Tong et al., 2024; Fatras et al., 2021). This gives a categorical Benamou-Brenier-type theorem for conditional flows with convex interpolants, the latter being the categorical analogue of shortest-path conditional continuous flows. When the similarity measure in the dynamic formulation is the discrete metric, the cost function in the Kantorovich formulation becomes the Hamming distance. When the similarity measure is the $L_{2}$ norm, the cost function is also $L_{2}$, mirroring the continuous case.
@@
An additional implication of stochastic crossing paths in DFM is that we cannot use an equivalent of the instantaneous change of variable formula (Chen et al., 2018) for probability estimation. Thus, other approaches are needed for estimating the perplexity. Inspired by the discrete diffusion bounds in Lou et al. (2024); Haxholli et al. (2025), we derive two upper bounds on perplexity for discrete flow matching. These bounds enable theoretically grounded training, model evaluation and comparison with other methods.
@@
Experiments show that minibatch-OT significantly reduces jumps in small-scale experiments and, in realistic settings (GPT2-sized model on OWT), reduces inference steps up to 32-fold (from 1024 to 32) to achieve the same generative perplexity. We also introduce multimask flows (DFM-MMF or simply MMF), which outperforms masked DFM in terms of generative perplexity, without sacrificing diversity, in particular when combined with minibatch-OT. Finally, we demonstrate that our derived bounds enable comparisons with autoregressive and discrete diffusion models.
@@
  • Finally, we introduce multimask flow (DFM-MM), which surpasses masked DFM models in generative perplexity without compromising diversity, with further gains achieved when applying OT. We show minibatch OT reduces inference steps up to 32-fold (1024 to 32) while maintaining performance in terms of generative perplexity on GPT2-scale models.
@@ 2 Preliminaries and Notation
A summary of Discrete Flow Matching is provided below. While the following preliminary is self-contained, we also provide a discrete diffusion introduction in Appendix D.
@@ 2.1.2 Conditional Probability Flows
where
$p_{t}^{i}(x^{i}|x_{0},x_{1})=(1-k_{t})\delta_{x_{0}^{i}}(x^{i})+k_{t}\delta_{x_{1}^{i}}(x^{i})$ (4)
with $k_{0}=0,k_{1}=1$ and increasing $k_{t}$. It is clear that this definition of $p_{t}(x|x_{0},x_{1})$ satisfies the conditions $p_{0}(x|x_{0},x_{1})=\delta_{x_{0}}(x)$ and $p_{1}(x|x_{0},x_{1})=\delta_{x_{1}}(x)$. In addition, Gat et al. (2024) show that component $i$ of the conditional probability velocity $u_{t}(x,z|x_{0},x_{1})$ corresponding to the flow defined in Equations (3) and (4) is
@@
Furthermore, they show that the probability velocity corresponding to the unconditional flow $p_{t}(z)$ can be written as a weighted sum of conditional probability velocities,
$u_{t}^{i}(x^{i},z)=\sum_{x_{0},x_{1}\in\mathcal{D}}u_{t}^{i}(x^{i},z|x_{0},x_{1})p(x_{0},x_{1}|z),$ (6)
which in the case of Equations (4) and (5) implies, $u_{t}^{i}(x^{i},z)=\frac{\dot{k}_{t}}{1-k_{t}}\left[p_{1|t}^{i}(x^{i}|z)-\delta_{z}^{i}(x^{i})\right].$ One then approximates $u_{t}^{i}(x^{i},z)$ by simply modeling $p^{i}_{1|t}(x^{i}|z)$ with a neural network $p^{i}_{1|t}(x^{i}|z;\theta)$ using the cross entropy loss $L$,
$\mathbb{E}_{t\sim U(0,1)}\mathbb{E}_{x_{0},x_{1}\sim\pi(x_{0},x_{1})}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0},x_{1})}\sum_{i=1}^{L}l_{t}^{i}(x_{1},x_{t},\theta),$ (7)
where $l_{t}^{i}(x_{1},x_{t},\theta)=-\log p^{i}_{1|t}(x_{1}^{i}|x_{t};\theta)$. It should be mentioned that in Gat et al. (2024), the definition of $p_{t}^{i}(x^{i}|x_{0},x_{1})$ is given in a more general form, but here we focus on this specific case for the sake of simplicity and since this formulation corresponds to shortest path conditional flows in the continuous framework, that is $X_{t}=(1-t)X_{0}+tX_{1}$.
@@ 3 Transition Reduction Objectives in Discrete Flow Matching
We first introduce our dynamic optimal transport formulation for discrete flows with per-position convex interpolants and derive its equivalent static (Kantorovich) formulation. Then, since masked flows (Gat et al., 2024) achieve the best practical performance but cannot leverage OT, we introduce multimasked flows. The latter maintain the performance and time-independent predictive probabilities of masked flows while enabling the application of OT approaches.
@@ 3.1 Dynamic OT Formulation of Discrete Flows with Convex Interpolants and Its Static Equivalent
with endpoints fixed at the source and target distributions. This objective is closely related to minimizing expected path length, but the squared speed penalizes large velocities more strongly. Moreover, by the Benamou–Brenier theorem (Benamou and Brenier, 2000; Tong et al., 2024), the infimum of Equation (8) equals the infimum of the Kantorovich (static) transport (functional) with quadratic cost,
$\int c(x_{0},x_{1})\pi(x_{0},x_{1})dx_{0}dx_{1},\text{ for }c(x_{0},x_{1})=\|x_{0}-x_{1}\|^{2}$ (9)
@@
$\int_{0}^{1}\sum_{x_{t}}\frac{1}{2}p(x_{t})\left[\sum_{i=1}^{L}\left(\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})-u_{t}^{i}(x_{t}^{i},x_{t})\right)\right]dt$
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})dt,$ (10)
@@
For these reasons, we define the categorical dynamic objective (functional) more generally as follows:
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\left[\sum_{i=1}^{L}\sum_{x^{i}}u_{t}^{i}(x^{i},x_{t})s(x^{i},x_{t}^{i})\right]dt$
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\left[\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})s(x^{i},x_{t}^{i})\right]dt,$ (11)
@@
Let $\pi(x_{0},x_{1})$ be the joint distribution of $x_{0}$ and $x_{1}$, and let $p_{t}$ be a flow defined as in Equations (2, 3, 4) that transforms $p=\sum_{x_{1}}\pi(x_{0},x_{1})$ into $q=\sum_{x_{0}}\pi(x_{0},x_{1})$. In this setting, the dynamic formulation given in Equation (11) equals the Kantorovich formulation:
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})s(x^{i},x_{t}^{i})dt$
$\sum_{x_{0},x_{1}}c(x_{0},x_{1})\pi(x_{0},x_{1}),$ (12)
@@
Theorem 3.1 addresses per-position discrete flows with convex interpolants, covering standard flows in the literature. By focusing on this specific setting and using path length instead of kinetic energy, our theorem unlike the classical Benamou–Brenier result, admits arbitrary similarity functions beyond quadratic costs. In addition, since the coupling $\pi$ fully determines the dynamics, we optimize only over $\pi$. Notably, dynamic and static losses match for every $\pi$ individually, meaning that the static and dynamic functionals are one and the same.
@@ 3.2 Multi-mask Flows
Standard masked DFM (Gat et al., 2024) uses a Dirac distribution at the fully masked sequence as its source, admitting only the trivial coupling. To enable meaningful couplings, we introduce multimask flow (DFM-MMF) with $V_{s}$ special mask tokens $m_{1},m_{2},...,m_{V_{s}}$}, all distinct from data tokens. Source sequences are uniformly sampled combinations of these masks. Our total vocabulary thus contains $V=V_{s}+V_{d}$ tokens: $V_{d}=50,257$ data tokens $x_{1},...,x_{V_{d}}$} from our tokenizer and $V_{s}$ mask tokens, i.e., $\mathcal{V}=\{x_{1},...,x_{V_{d}},m_{1},m...,m_{V_{s}}\}$. At $t=0$, only mask tokens appear while data tokens are assigned zero probability. This design provides two advantages: denoising probabilities remain time-independent as in masked flow (see Appendix A.6), and mask embeddings are fully unrestricted since they are untied from data-token embeddings. Effectively, we create a "fictitious grid" where each $L$-length sequence holds mass $\varepsilon=\frac{1}{V_{s}^{L}}$ for large enough $V_{s}$. The flow transports these small masses to the data grid, enabling OT.
@@ 4 Upper Bounds on the Perplexity in Discrete Flow Matching
Perplexity is a key evaluation metric for language models, making it essential to calculate or bound it in the DFM framework. While we provide a precise but computationally challenging formula in Appendix A.5, the next two subsections present practical bounds. These bounds function as both principled training objectives and practical evaluation metrics for general discrete flows, offering the same intrinsic and objective assessment as discrete diffusion bounds.